Please help me solve these questions:
1) (3x-6)/(x^2-4x+4) Reduce to lowest terms
2) (x)/(3)= (4)/(x+1) Solve the equation. 2 answers
3) (-3)/(x) = (x)/(-3) Solve the equation. 2 answers

Answer by checkley77(12844)

1) (3x-6)/(x^2-4x+4) Reduce to lowest terms
3(X-2)/(X-2)^2 CANCEL OUT THE (X-2) TERMS.
3/(X-2) ANS.
2) (x)/(3)= (4)/(x+1) Solve the equation. 2 answers
X(X+1)=3*4
X^2+X=12
X^1+X-12=0
(X+4)(X-3)=0
X+4=0
X=-4 ANS
X-3=0
X=3 ANS.
3) (-3)/(x) = (x)/(-3) Solve the equation. 2 answers
X^2=9
X^2-9=0
(X-3)^2=0
X-3=0
X=3 BOTH ANS.